OpenStudy (anonymous):

If i apply force to an object, but the object is too heavy to have moved, will i still have done work?

OpenStudy (mayankdevnani):

yaa

OpenStudy (mayankdevnani):

@aceace

OpenStudy (mayankdevnani):

confusion?? because w=fs

OpenStudy (mayankdevnani):

force is applied and displacement is also occur

OpenStudy (mayankdevnani):

@aceace

OpenStudy (anonymous):

but there is no displacementas there is too much mass then will there have been work done?

OpenStudy (amistre64):

it depends on where your definition of "work" is focused.

OpenStudy (anonymous):

no movement, no work done... \[\int\limits_{a}^{a} F*dx = 0 \] ...no matter what F is

OpenStudy (anonymous):

The phrase "i did work" still unclear. Only force do work. If you refer word ""I" to the "force you used to push the wall", then work is zero. But there's another force when you push the wall that does work. That's why you feel tired even if the wall isn't moving :)

OpenStudy (anonymous):

work done = force * displacement so if object is not moving then you have done no work done

OpenStudy (anonymous):

no displacement in the direction of force, no work done. even if you died pushing the object hard. ^_^ No displacement along the direction of force means no work done.

OpenStudy (anonymous):

w=\[\int\limits_{0}^{d}\]f.ds...hence if displacement is zero,no work is said to be done in physics...